Give an example of a divergent sequence which has two convergent sequences. Justify  your claim


1
Expert's answer
2021-04-14T13:17:24-0400

"\displaystyle
\textsf{Let}\,\, a_n = \{(-1)^n\}_{n = 1}^{\infty} = \{-1, 1, -1, 1, -1, 1,\ldots\} \\

\textsf{Let}\,\, a_{n1}\,\, \textsf{and}\,\, a_{n2}\,\, \textsf{be subsequences of}\,\, a_n \\

\textsf{Let}\,\,\, a_{n1} = \{-1,-1,-1,\ldots\} \\

a_{n1} \,\,\, \textsf{converges to}\,\, -1 \\

\textsf{Let}\,\,\, a_{n2} = \{1,1,1,\ldots\} \\

a_{n2} \,\,\, \textsf{converges to}\,\, 1 \\
\textbf{Proof that}\,\,\textbf{a}_n \,\, \textbf{diverges} \\

\textsf{Assume that}\,\, a_n\,\, \textsf{converges to}\,\, a^*\,\, \textsf{for any}\,\, \epsilon > 0.\\

|a_n - a^*| < \epsilon\\

\textsf{Let}\,\, \epsilon = \frac{1}{2} \\
|(-1)^n - a^*| < \frac{1}{2}\\

|(-1)^{n + 1} - (-1)^{n + 2}| = 2 \\

|(-1)^{n + 1} - a^* + a^* - (-1)^{n + 2}| = 2 \\

\begin{aligned}
2 &= |(-1)^{n + 1} - a^* + a^* - (-1)^{n + 2}| 
\\&\leq |(-1)^{n + 1} - a^*| + |a^* - (-1)^{n + 2}| \\
&\leq |(-1)^{n + 1} - a^*| + |(-1)^{n + 2} - a^*| < \frac{1}{2} + \frac{1}{2} = 1 
\end{aligned}\\

\textsf{This implies that}\,\, 2 <1, \,\, \textsf{which is false}.\\

\textsf{Hence the sequence is divergent.}"